# Large Language Models for Information Retrieval: A Survey

### v4

@@ 5.4 Limitations
Although recent research on utilizing LLMs for document reranking has made significant progress, it still has some limitations. First, due to the reliance on API calls and a large number of parameters, the process of LLM ranking could be expensive and inefficient. Therefore, achieving a trade-off between the cost/efficiency and performance of LLMs is a topic worth discussing. Along this line,  Rashid et al. [179] propose a budget-aware ranking solution which maximizes the LLM’s performance within a given budget. Besides,  Meng et al. [180] systematically discuss the improvements in ranking efficiency and effectiveness brought by the rank list truncation technique. Second, while existing studies mainly focus on applying LLMs to open-domain datasets (such as MSMARCO [99]) or relevance-based text ranking tasks, their adaptability to in-domain datasets [113] and non-standard ranking datasets [181] remains an area that demands more comprehensive exploration.
@@ 6.1 Passive Reader
To generate answers for users, a straightforward strategy is to supply the retrieved documents according to the queries or previously generated texts from IR systems as inputs to LLMs for creating passages [194, 182, 184, 185, 186, 187, 195, 196, 191, 183, 189, 23, 192, 197]. By this means, these approaches use the LLMs and IR systems separately, with LLMs functioning as passive recipients of documents from the IR systems. The strategies for utilizing LLMs within IR systems’ reader modules can be categorized into the following three groups according to the frequency of retrieving documents for LLMs.
@@ 7 Search Agent
With the development of LLMs, IR systems are also facing new changes. Among them, developing LLMs as intelligent agents has attracted more and more attention. This conceptual shift aims to mimic human browsing patterns, thereby enhancing the capability of these models to handle complex retrieval tasks. Empowered by the advanced natural language understanding and generation capabilities of LLMs, these agents can autonomously search, interpret, and synthesize information from a wide range of sources.
One way to achieve this ability is to design a pipeline that combines a series of modules and assigns different roles to them. Such a pre-defined pipeline mimics users’ behaviors on the web by breaking it into several sub-tasks which are performed by different modules. However, this kind of static agent cannot deal with the complex nature of users’ behavior sequences on the web and may face challenges when interacting with real-world environments. An alternative solution is to allow LLMs to freely explore the web and make interactions themselves, namely letting the LLM itself decide what action it will take next based on the feedback from the environment (or humans). These agents have more flexibility and act more like human beings.
7.1 Static Agent
To mimic human search patterns, a straightforward approach is to design a static system to browse the web and synthesize information step by step [258, 259, 260, 261, 262, 263]. By breaking the information-seeking process into multiple subtasks, they design a pipeline that contains various LLM-based modules in advance and assigns different subtasks to them.
LaMDA [258] serves as an early work of the static agent. It consists of a family of Transformer-based neural language models specialized for dialog, with up to 137B parameters, pre-trained on 1.56T tokens from public dialogue data and web text. The study emphasizes the model’s development through a static pipeline, encompassing large-scale pre-training, followed by strategic fine-tuning stages aimed at enhancing three critical aspects: dialogue quality, safety, and groundedness. It can integrate external IR systems for factual grounding. This integration allows LaMDA to access and use external and authoritative sources when generating responses. SeeKeR [259] also incorporates the Internet search into its modular architecture for generating more factual responses. It performs three sequential tasks: generating a search query, generating knowledge from search results, and generating a final response. GopherCite [262] uses a search engine like Google Search to find relevant sources. It then synthesizes a response that includes verbatim quotes from these sources as evidence, aligning the Gopher’s output with verified information. WebAgent [261] develops a series of tasks, including instruction decomposition and planning, action programming, and HTML summarization. It can navigate the web, understand and synthesize information from multiple sources, and execute web-based tasks, effectively functioning as an advanced search and interaction agent. WebGLM [260] designs an LLM-augmented retriever, a bootstrapped generator, and a human preference-aware scorer. These components work together to provide accurate web-enhanced question-answering capabilities that are sensitive to human preferences. Shi et al. [263] focus on enhancing the relevance, responsibility, and trustworthiness of LLMs in web search applications via an intent-aware generator, an evidence-sensitive validator, and a multi-strategy supported optimizer. CoSearchAgent [264] leverages LLMs to understand and process user queries within the context of multi-party conversations. It is capable of rewriting incomplete or ambiguous queries, asking clarifying questions, and searching the web for relevant information using APIs.
7.2 Dynamic Agent
Researches of dynamic agent can be further categorized into agent design, framework and system, and benchmark.
7.2.1 Agent Design
Instead of statically arranging LLMs in a pipeline, WebGPT [24] takes an alternate approach by training LLMs to use search engines automatically. This is achieved through the application of a reinforcement learning framework, within which a simulated environment is constructed for GPT-3 models. Specifically, the WebGPT model employs special tokens to execute actions such as querying, scrolling through rankings, and quoting references on search engines. This innovative approach allows the GPT-3 model to use search engines for text generation, enhancing the reliability and real-time capability of the generated texts. A following study [265] has extended this paradigm to the domain of Chinese question answering. ASH (Actor-Summarizer-Hierarchical) prompting [266] significantly enhances the ability of LLMs on WebShop benchmark. TRAD [267] enhances LLM agents for sequential decision-making tasks. It first utilizes step-level demonstration selection through thought matching. Then it enhances the agent’s decision-making process by considering the temporal context of actions. AutoWebGLM [268] aims address the limitations of current web browsing agents, which often struggle with the versatility of webpage actions, the length of HTML text, and the open-domain nature of web decision-making. It uses an HTML simplification algorithm, a hybrid human-AI method and reinforcement learning and rejection sampling to bootstrap the model, allowing it to improve webpage comprehension, browser operations, and task decomposition autonomously.
7.2.2 Framework and System
KwaiAgents [269] is a generalized information-seeking agent system built around LLMs. In this system, LLMs act as the cognitive core, capable of understanding queries, referencing external documents, updating and retrieving internal memory, and executing tasks using a hybrid search-browse toolkit. WebVoyager [270] is a large multimodal model-powered agent, capable of navigating and interacting with websites by processing both visual inputs (like screenshots) and textual content, mimicking human web browsing behavior. IoA [271] is a framework that enables collaboration among a diverse range of autonomous agents, particularly those based on LLMs. It aims to address the issue of integrating third-party agents, simulating distributed environments, and adapting to dynamic task requirements due to rigid communication and coordination mechanisms. Agent-E [272] aims to address the problems of dynamic and noisy nature of web content, the complexity of user interfaces, and the need for sophisticated planning to execute multi-step tasks. It utilizes a two-tiered system consisting of a planner agent and a browser navigation agent, which allows for a clear separation of roles and responsibilities. Besides, Agent-E introduces a flexible approach to Document Object Model (DOM) distillation, enabling it to choose the most suitable DOM representation for a given task. MindSearch [273] is a multi-agent framework that mimics human cognitive processes in web information seeking and integration. It involves a two-pronged approach using a WebPlanner and multiple WebSearchers. The WebPlanner acts as a high-level planner, decomposing user queries into atomic sub-questions and orchestrating the reasoning process. The WebSearchers, are tasked with performing hierarchical information retrieval from the internet based on the sub-questions assigned by the WebPlanner. They employ a coarse-to-fine selection strategy to efficiently extract valuable information from a large number of web pages, managing the information overload that LLMs might otherwise struggle with.
7.2.3 Benchmark
Some works develop important benchmarks for interactive web-based agents [274, 275, 276, 277]. For example, WebShop [275] aims to provide a scalable, interactive web-based environment for language understanding and decision-making, focusing on the task of online shopping. It first takes a raw observation from the environment and produces a new, more meaningful representation that aligns with the specific goal. Then, it dynamically predicts the next action based on the summarized observation and the interaction history. WebCanvas [277] is also an online evaluation framework designed to benchmark the performance of web agents in dynamic and evolving online environments. It aims to address the gap between existing benchmarks, which tend to capture static aspects of the web, and the real-world needs of web agents that must adapt to frequent updates in user interfaces and content.
7.3 Limitations
Though the aspect of static search agents has been thoroughly studied, the literature on dynamic search agents remains limited. Some agents may lack mechanisms for real-time fact-checking or verification against authoritative sources, leading to the potential dissemination of misinformation. Moreover, since LLMs are trained on data from the Internet, they may inadvertently perpetuate biases present in the training data. This can lead to biased or offensive outputs and may collect unethical content from the web. Finally, as LLMs process user queries, there are concerns regarding user privacy and data security, especially if sensitive or personal information is involved in the queries.

### v5

@@ 5.5 Limitations
Although recent research on utilizing LLMs for document reranking has made significant progress, it still has some limitations. First, due to the reliance on API calls and a large number of parameters, the process of LLM ranking could be expensive and inefficient. Therefore, achieving a trade-off between the cost/efficiency and performance of LLMs is a topic worth discussing. Along this line,  Rashid et al. [198] propose a budget-aware ranking solution which maximizes the LLM’s performance within a given budget. Notably, Chen et al. [199] introduce in-context re-ranking (ICR), an attention-based method that achieves superior efficiency by eliminating generative overhead through O(1) forward passes. Besides,  Meng et al. [200] systematically discuss the improvements in ranking efficiency and effectiveness brought by the rank list truncation technique. Second, while existing studies mainly focus on applying LLMs to open-domain datasets (such as MSMARCO [102]) or relevance-based text ranking tasks, their adaptability to in-domain datasets [120], non-standard ranking datasets [201] and reasoning-intensive datasets [196] remains an area that demands more comprehensive exploration.
@@ 6.1 Passive Reader
To generate answers for users, a straightforward strategy is to supply the retrieved documents according to the queries or previously generated texts from IR systems as inputs to LLMs for creating passages [214, 202, 204, 205, 206, 207, 215, 216, 211, 203, 209, 23, 212, 217, 218]. By this means, these approaches use the LLMs and IR systems separately, with LLMs functioning as passive recipients of documents from the IR systems. The strategies for utilizing LLMs within IR systems’ reader modules can be categorized into the following three groups according to the frequency of retrieving documents for LLMs.
@@ 7 Search Agent
The emergence of large reasoning models (LRMs) has ushered in a new era for IR systems, with a growing focus on developing LRM-based intelligent agents. This paradigm shift seeks to replicate human-like reasoning and retrieval processes, thereby augmenting the capacity of LLM-powered IR models to tackle complex, real-world problems. Leveraging their advanced natural language understanding, reasoning, and generation capabilities, these agents can autonomously search, interpret, and synthesize information from diverse sources.
Initial research in this domain focused on static pipeline-based architectures, where an information-seeking task is broken down into a series of modules, each with a pre-defined role [279, 280, 281, 282, 283, 284, 285]. While these systems demonstrate a foundational approach, their fixed workflows limit their ability to adapt to the dynamic and complex interactions inherent in real-world scenarios. This inflexibility constrains their overall performance and hinders their effectiveness in advanced reasoning and problem-solving.
Recently, the development of LRMs has enabled the development of a new class of autonomous search agents. These agents move beyond static pipelines by allowing the LLM to actively and dynamically explore the web. This is achieved by enabling the model to decide its next action based on real-time feedback from the environment or humans. This shift towards flexible, self-guided behavior makes these agents more adaptable and more closely aligned with human-like problem-solving.
In this section, we will comprehensively introduce the studies about search agents from the following four aspects: (1) architecture of search agents, (2) information seeking module, (3) optimization of search agents, and (4) benchmarks and resources.
7.1 Architecture of Search Agent
The design of a search agent’s architecture is a foundational step that establishes its core operational mechanism. Existing approaches can be broadly categorized into two main paradigms: single-agent frameworks and multi-agent frameworks. A single-agent framework utilizes a single LLM to handle all aspects of the task, including reasoning, interaction, and answer generation. In contrast, a multi-agent framework distributes these responsibilities among multiple LLMs, which act as collaborative agents to achieve the target.
7.1.1 Single-Agent Frameworks
In a single-agent framework, a single LLM with strong reasoning capabilities serves as the central decision-maker and task executor. This LLM dynamically determines the next action based on the current information state and environmental context, managing the entire reasoning and interaction process. For example, Search-R1 [286] and ReSearch [287] adopt a ReAct-style mechanism, enabling the LLM policy to automatically generate actions such as “think”, “search”, and “answer”. This allows the agent to iteratively interact with search tools to resolve complex multi-hop questions. To optimize the performance of these single-agent systems, researchers have employed various reinforcement learning (RL) techniques. The GRPO algorithm [288] has been commonly utilized to enhance performance. In a more refined approach, R1-Searcher [289] introduces a two-stage GRPO-based RL optimization. The first stage assigns rewards based on retrieval frequency and output format correctness, while the second stage rewards answer accuracy and the final output format. To further improve reasoning and exploration, Li et al. [290] propose START, which introduces a hint-infer mechanism that manually inserts hint strings during inference. This encourages the LLM to self-reflect and make better use of external tools. They also design Hint-RFT, a method that performs rejection sampling and revises reasoning trajectories to support the supervised fine-tuning of search agents. More recently, Atom-Searcher [291] has been proposed to decompose the holistic “thinking” process into several finer-grained “atom-thinking” actions. These actions are guided by reasoning reward models, which provide precise feedback on reasoning trajectories, going beyond simple outcome-based rewards.
The main advantage of the single-agent framework lies in its simplicity, as it can be trained end-to-end via RL. This allows researchers to explore the reasoning limits of a single model using carefully designed optimization algorithms. However, a single agent often struggles with highly complex queries that require extensive tool use and long-context reasoning. To address this limitation, multi-agent frameworks have been explored, where multiple agents collaborate to complete complex search and reasoning tasks.
7.1.2 Multi-Agent Frameworks
Multi-agent frameworks utilize multiple, specialized LLMs that collaborate to complete complex search and reasoning tasks. This approach enables a division of labor, where each agent focuses on a distinct function, leading to improved system effectiveness and efficiency. For example, KwaiAgents [292] separates reasoning and summarization into two distinct agents. The reasoning agent is equipped with capabilities such as query understanding, external documents referencing, memory management, and task execution via a hybrid search–browse toolkit. Similarly, Mei et al. [293] propose decoupling the search agent into a search planner and a generator, with optimization efforts primarily on the planner. Building on this, MindSearch [294] introduces a multi-agent framework inspired by human cognitive processes for web retrieval. Its architecture consists of a WebPlanner and multiple WebSearchers. The WebPlanner acts as a high-level controller, decomposing user queries into atomic sub-questions and designing the reasoning process. The WebSearchers then perform hierarchical retrieval over the web, guided by these sub-questions. By employing a coarse-to-fine selection strategy, they efficiently filter valuable information from a large pool of web pages, thereby alleviating the information overload that often hinders LLMs. Alita [295] advances this idea by incorporating self-evolving capabilities through a dynamic MCP box. Its manager agent handles central task planning and MCP brainstorming, deciding whether to generate new MCP tools for emerging tasks. The web agent, in turn, is responsible for browsing and retrieving external information. Similarly, OWL [296] further decouples central planning and task execution to improve generalization across different domains. OWL consists of a domain-agnostic planner agent for high-level task decomposition, a coordinator agent for managing task assignments and dependencies, and a set of specialized agents with domain-specific toolkits that execute subtasks and report results. This modular design allows researchers to focus optimization efforts on the planner, enhancing its adaptability while minimizing training complexity for other components.
The primary advantage of the multi-agent framework is that it allows individual agents to specialize in distinct tasks, enhancing overall system effectiveness and efficiency. However, it introduces challenges in jointly optimizing multiple agents through RL. Current approaches often limit optimization to the core planner agent, which plays the most critical role in coordinating the framework. Therefore, developing more stable and efficient RL strategies for multi-agent search systems remains a promising direction for future research.
7.2 Information Seeking Module
To handle atomic queries or complete sub-tasks from an upstream planner, a search agent typically relies on an information-seeking module that locates, collects, and synthesizes relevant information. These approaches can be roughly categorized into two types: API-based and browsing-based methods. API-based methods use search engine APIs to retrieve information, while browsing-based methods construct sandboxes or virtual environments that enable agents to simulate human-like web interactions.
7.2.1 API-based Information Seeking
The most straightforward information-seeking strategy is to leverage search engine and scientific database APIs as external tools. Many commercial applications, such as Gemini DeepResearch and Grok DeepSearch, [^7] Gemini DeepResearch: https://gemini.google/overview/deep-research/, Grok DeepSearch: https://x.ai/news/grok-3 rely on APIs like Google Search, Bing Search, and X Search to access external knowledge. In the research community, Cognitive Kernel-Pro [297] uses the free DuckDuckGo search interface to create a fully open-source pipeline, while CoSearch-Agent [298] integrates SerpApi for real-time search within Slack-based environments. Beyond basic search APIs, other systems incorporate specialized APIs to refine the retrieval process. For example, Search-o1 [299] and Agent Laboratory [300] use Jina Reader API to extract and refine web passages for downstream reasoning, and the arXiv API to obtain academic metadata. Similarly, AI Scientist [301] employs the Semantic Scholar API to verify citation relationships. While simple and accessible, API-based approaches often struggle with complex, dynamic content rendered by JavaScript, interactive components, or information gated behind authentication.
7.2.2 Browsing-based Information Seeking
In contrast to API-based methods, browsing-based information-seeking approaches provide search agents with interactive environments that simulate human-web interactions. For example, Manus AI’s browsing agent creates a sandboxed Chromium instance for each subtask, [^8] Manus AI: https://manus.im/ while AutoGLM  [^9] AutoGLM: https://autoglm-research.zhipuai.cn/ sequentially opens web pages, reads content, and generates refined reports. In research, AutoAgent [302] uses the BrowserGym environment to perform scrolling and interaction with webpage components. SimpleDeepSearcher [303] and Tool-star [304] further compress retrieved content from both browsing and API-based extraction to generate condensed references for answer generation. While browsing-based approaches are better suited for retrieving real-time and deeply nested content, they typically incur higher latency and resource costs. To address the high costs associated with real search APIs, some recent works explore alternative approaches. For example, ZeroSearch [305] trains LLMs to simulate search engine behavior without making actual API calls, thereby significantly reducing training costs. Additionally, some methods, such as Alita [295], propose to dynamically create new MCP tools during the agent’s reasoning process, enabling a self-evolving capability that reduces reliance on pre-defined toolkits and further optimizes computation costs.
7.3 Optimization
To transform general-purpose LLMs into specialized search agents, researchers have explored various optimization and fine-tuning methods. These approaches aim to internalize advanced search skills, such as planning, reasoning, and tool usage into the model’s parametric knowledge. The ultimate goal is to enable agents to perform exploratory information acquisition. Based on the progressive levels of search capabilities, this section categorizes mainstream agent tuning methods into three paradigms: strategic retrieval optimization, iterative search tuning, and autonomous open-web search.
7.3.1 Strategic Retrieval Optimization
In basic search scenarios, agents must first learn when to retrieve information and how to formulate high-quality queries. Early RAG methods typically follow a fixed and passive pipeline: given a question, the system directly performs a search and then generates an answer based on the results. This approach is often inefficient, as it can lead to redundant searches and struggles to handle irrelevant information.
Recent studies have introduced strategic retrieval optimization techniques that enable agents to explicitly model retrieval decisions, thereby balancing search costs against the potential benefit of new information. For example, Open-RAG [306] proposes a “hybrid adaptive retrieval” mechanism that learns to generate specific tokens to control its retrieval behavior. This work also introduces a constructive learning paradigm, which actively injects distracting information into training data to enhance the model’s robustness and discrimination capabilities against noisy or irrelevant search results. Similarly, DeepRAG [307] models retrieval decisions as a Markov decision process, using imitation learning to train models to weigh the benefits of relying on internal knowledge versus performing an external search at each reasoning step. This provide the model with the dynamic ability to decide when to retrieve. ATLAS [308] takes a different approach by applying gradient backpropagation only on “critical steps” within expert trajectories. In search tasks, these steps correspond to key decisions such as initiating a search or formulating a core query. By focusing the training signal on these strategic points, this method improves the agent’s core decision-making and generalization ability.
7.3.2 Iterative Retrieval Tuning
For complex tasks that cannot be solved with a single retrieval, search agents require multi-step reasoning and iterative information acquisition. These capabilities are crucial for applications such as multi-hop question answering and open-domain problem solving, where agents must converge on an answer through a dynamic cycle of “think-search-integrate-rethink”. Research in this area has leveraged both supervised and reinforcement learning paradigms.
Supervised learning trains models by providing expert trajectories that include all intermediate reasoning and retrieval steps. CoRAG [309] exemplifies this approach by automatically generating retrieval chains with intermediate sub-queries and sub-answers for existing datasets. Through rejection sampling, it enables models to explicitly learn multi-step retrieval patterns. Similarly, Auto-RAG [310] focuses on synthesizing instruction data that contains retrieval decision processes, allowing models to master autonomous multi-round retrieval logic through SFT.
Reinforcement learning allows agents to autonomously explore and learn optimal dynamic search strategies through interaction with an environment. Works such as ReSearch [287], R1-Searcher [289], and Search-R1 [286] all adopt RL frameworks, defining search as a learnable action within the reasoning process. Agents learn when and how to intersperse search queries by maximizing rewards from task success. Among these, R1-Searcher designs a two-stage RL training pipeline that effectively decouples the objectives of learning to use tools from learning to solve problems with tools. At the algorithmic level, ARPO [311] optimizes training efficiency for iterative agents by proposing an entropy-based adaptive exploration mechanism. This method increases exploration intensity at critical decision points where models show high uncertainty, significantly reducing training costs.
7.3.3 Autonomous Open-Web Search
At the highest level, search agents must be able to operate autonomously within open and dynamic web environments. This capability requires models to handle complex challenges such as noisy data, conflicting information from multiple sources, and a lack of explicit supervision. To succeed, they must possess advanced skills in information planning, cross-validation, and multimodal understanding.
Recent studies show that end-to-end RL is an effective path for endowing models with autonomous research capabilities. WebAgent-R1 [312] and DeepResearcher [313] use sparse reward training in real browser environments, enabling agents to autonomously plan search paths, verify information, and integrate knowledge from various sources. WebThinker [314] proposes an “Think-Search-and-Draft” strategy, using iterative online direct preference optimization (DPO) to enable agents to seamlessly switch between information collection, reasoning, and content generation.
Subsequent research has focused on building more systematic training methodologies. Works such as WebDancer [315], WebSailor [316], and WebShaper [317] demonstrate that a combination of high-quality data synthesis and hybrid training strategies is an effective path for training advanced agents. These works have made important innovations at the data level. For example, WebShaper proposes a “formalization-driven” data synthesis framework that generates logically consistent data from a task’s reasoning structure. WebWatcher [318] further advances this filed by incorporating visual information during training, enabling models to understand and utilize both images and texts on web pages, thereby moving toward human-like research capabilities. Through these methods, search agents are gradually evolving into research-oriented agents capable of autonomous exploration and information integration on the open web.
7.4 Benchmarks and Resources
To effectively evaluate and advance search agents, a diverse set of benchmarks and resources is essential. Current evaluation methodologies can be broadly categorized into two main paradigms: QA-style benchmarks, which assess the agent’s ability to answer complex questions, and task-oriented benchmarks, which measure its capacity for planning, tool use, and environmental interaction. Additionally, a growing number of agent platforms and datasets serve as valuable resources for both evaluation and model training.
7.4.1 QA Benchmarks
QA benchmarks are designed to evaluate problem-solving and reasoning capabilities of search agents. They range from simple factual recall to multi-hop reasoning and expert-level challenges.
Single-hop QA. It involves questions that can be answered by retrieving information from a single document or source. These benchmarks, such as TriviaQA [319], SimpleQA [320], PopQA [321], and Natural Questions (NQ) [103], primarily test a model’s ability to perform open-domain factual retrieval and reading comprehension. For example, NQ provides anonymized Google search queries paired with human-annotated answers and Wikipedia evidence.
Multi-hop QA. It requires models to reason over and combine information from multiple sources to find the answer. A typical multi-hop QA dataset is HotpotQA [322]. It focuses on multi-hop reasoning by providing questions that require chaining evidence across multiple Wikipedia pages. 2WikiMultiHopQA [323] extends this by mixing structured knowledge and unstructured text and providing explicit reasoning paths for a more fine-grained evaluation of multi-step inference.
Expert-level Challenges. They are designed to be extremely difficult, often requiring deep domain knowledge and complex reasoning to push the limits of advanced models. Humanity’s Last Exam (HLE) [324] assembles thousands of hard, expert-crafted questions to stress test models across broad domains. Benchmarks like BrowseComp [325] attempt to force genuine web-based retrieval by filtering out items solvable from parametric memory. However, even with these efforts, top-performing systems can still exploit internal knowledge, which may overstate their true research capability.
7.4.2 Task-oriented Benchmarks
Task-oriented benchmarks assess an agent’s practical skills in planning, tool use, and interaction with various environments.
General Assistant Workflows. GAIA [326], AssistantBench [327], and Magnetic-One [328] cover broad assistant tasks that require planning across dialogue, retrieval, and simple tool calls. GAIA, for instance, measures an agent’s end-to-end task management, while Magnetic-One emphasizes robustness across diverse domains and chained subtasks.
Code and Research. SWE-bench [329], HumanEvalFix [330], MLE-bench [331], and MLAgentBench [332] probe pipelines centered on software engineering and research. They require agents to perform tasks like code implementation, debugging, experiment setup, and hyperparameter tuning.
Multi-Agent Coordination. RE-Bench [333] and RESEARCHTOWN [334] stress multi-agent collaboration, role assignment, and iterative refinement on shared research goals.
GUI Control. WebArena [335] and SpaBench [336] extend evaluation to include direct interface manipulation, measuring an agent’s ability to control web UIs or simulated devices and handle noisy, stateful environments.
7.4.3 Agent Datasets and Platforms
Beyond static benchmarks, several projects offer comprehensive resources that bundle evaluation suites with data-generation pipelines and agent toolkits. These resources serve as both benchmarks and valuable training corpora for agent research.
The Alibaba-NLP WebAgent repository is a notable example, packaging a web-traversal benchmark (WebWalkerQA [337]) with agent models and data tools (WebDancer [315], WebShaper [317], and WebSailor [316]). Specifically, WebWalkerQA [337] probes an agent’s ability to traverse sites and extract evidence across multiple subpages, emphasizing structured navigation over single-turn retrieval. WebDancer [315] implements a four-stage training paradigm and releases both models and browsing trajectories, enabling reproducible, end-to-end evaluation. WebShaper [317] provides a “formalization-driven” data synthesis pipeline that systematically generates information-seeking instances, making it valuable for cold-starting agents and for studying data-centric training strategies. The recent model releases from WebSailor [316] demonstrate how post-training and specialized agent tuning can yield stronger navigation and planning behaviors on these benchmarks.